OpenStudy (anonymous):

Converting 3x^2 + 6y^2 = 18 to standard form of a eclispe

OpenStudy (kirbykirby):

Well the general form of an ellipse is... \[ \frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] However, you just have one "x^2" term and one "y^2" term which should lead you into seeing that you have the ellipse centered at the origin (h=0, k=0), so \[ \frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]. Your equation is very similar to this!! A big hint to get the \(=1\) part is to divide both sides by 18 .